# On Geometry and Topology of Hessian Manifolds

### v4

@@ 2.3.2 Applications of Koszul Forms
Since $\mathrm{H}^{1}(M;\mathbb{R})=0$, by Proposition 2.32 we have that the Picard group of $M$ is isomorphic to $\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})$, and hence is 2-torsion. In particular, the canonical line bundle $K$ of $(M,\nabla)$ satisfies $K^{\otimes 2}=0$. Since $M$ is orientable, the $\mathbb{R}$-line bundle $\bigwedge^{n}(T^{*}M)$ is trivial. Therefore, by virtue of Lemma 2.35, the canonical line bundle of $M$ is trivial as a flat $\mathbb{R}$-line bundle. Corollary 2.37 then implies that $M$ admits a volume form $\Omega$ such that $\nabla\Omega=0$. The desired result follows immediately from Theorem 2.44. ∎
As an application of the first Koszul forms, we suggest several important splitting theorems of Hessian manifolds, which all relies on the following fundamental result of Tischler on fibration over circles.
Theorem 2.46 (Tischler, 1970[19]).
Theorem 2.47.
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type. Then there exists a fiber bundle $\varphi\colon M\rightarrow\mathbb{S}^{1}$.
Proof.
Let $\eta\in\Omega^{1}(M)$ be a $1$-form potential of $((M,\nabla),h)$. View $\eta\colon M\rightarrow T^{*}M$ as an embedding. Then, $\eta$ vanishes at $p\in M$ if and only if $p$ is an intersection point of $\eta\colon M\rightarrow T^{*}M$ and the zero section $\sigma$ of the cotangent bundle of $M$. Since $\nabla\eta=h>0$, each intersection point of $\eta$ and $\sigma$ is a nondegenerate transverse intersection, and the intersection numbers at those intersection points are all positive. But, on the other hand, the sum of the intersection numbers at the intersection points of $\eta$ and $\sigma$ is precisely
$\langle e(T^{*}M),[M]\rangle=\chi(M)=0$
by Proposition 2.17. This proves that the sections $\eta$ and $\sigma$ must have no intersection, i.e. $\eta$ is a nowhere vanishing closed 1-form on $M$. The desired result then follows from Theorem 2.46. ∎
An stronger result can be obtained if the manifold under consideration is orientable.
Theorem 2.48.
Proof.
Let $\varkappa$ be the first Koszul form of $((M,\nabla),h)$. If $\varkappa=0$, then $(M,h)$ is a flat Riemannian manifold by Theorem 2.44. We therefore assume $\varkappa\neq 0$. But then since $\bar{\nabla}\varkappa=0$ by Proposition 2.43, we have that $\varkappa$ is nowhere vanishing by parallel transport.
Since $\mathrm{H}^{1}(M;\mathbb{Q})$ is dense in $\mathrm{H}^{1}(M;\mathbb{R})$, there exists a closed $1$-form $\hat{\varkappa}\in\Omega^{1}(M)$ such that $\hat{\varkappa}]\in\mathrm{H}^{1}(M;\mathbb{Q})$ and $\hat{\varkappa}$ is also nowhere vanishing. Define
$m:=\operatorname{min}\{k\in\mathbb{R}:k>0,k[\hat{\varkappa}]\in\mathrm{H}^{1}(M;\mathbb{Z})\}.$
Fix a base point $q$ on $M$, and define the smooth mapping $\varphi\colon M\rightarrow\mathbb{S}^{1}$ via
[…]
where $\gamma\colon[0,1]\rightarrow M$ is a smooth curve with $\gamma(0)=q$ and $\gamma(1)=p$. Since the periods of $m[\hat{\varkappa}$] are integers, the map $\varphi$ is well-defined. Moreover, since $\hat{\varkappa}$ is nowhere vanishing, we have that $\varphi$ is a submersion. By Ehresmann’s fibration theorem, we have that $\varphi\colon M\rightarrow\mathbb{S}^{1}$ is a fiber bundle. Denote by $F$ the fiber of $\varphi\colon M\rightarrow\mathbb{S}^{1}$.
Identifying each homotopy class in $\pi_{1}(\mathbb{S}^{1})$ with its winding number, it is readily seen that the induced homomorphism $\varphi_{*}\colon\pi_{1}(M)\rightarrow\pi_{1}(\mathbb{S}^{1})=\mathbb{Z}$ satisfies
$\varphi_{*}([\gamma])=\int_{\gamma}m\hat{\varkappa}$
for all $\gamma]\in\pi_{1}(M)$. Since $\hat{\varkappa}\neq 0$, we have that $\operatorname{Im}(\varphi_{*})\neq 0$. Therefore, there exists a positive integer $d\in\mathbb{Z}$ such that $\operatorname{Im}(\varphi_{*})=d\mathbb{Z}$. But by the minimality of the factor $m$, we must have that $d=1$ and hence $\operatorname{Im}(\varphi_{*})=\mathbb{Z}$. Now, since $M$ is connected and $\varphi_{*}$ is surjective, the long exact sequence
Moreover, since $\varphi\colon M\rightarrow\mathbb{S}^{1}$ is a submersion between compact differentiable manifolds, we have that $F$ is also a compact differentiable manifold.
By the very definition of $\varphi$, it is readily seen that the vector field $\hat{\varkappa}^{\sharp}$ is transverse to the fibers of $\varphi\colon M\rightarrow\mathbb{S}^{1}$. Consider the (n1)(n-1)-form $\Omega$ on $M$ defined by
$\Omega(X_{1},\dots,X_{n-1}):=\operatorname{vol}_{h}(X_{1},\dots,X_{n-1},\hat{\varkappa}^{\sharp})$
where $X_{1},\dots,X_{n-1}$ are arbitrary vector fields on $M$. Then by construction $\Omega|_{F}$ is nowhere vanishing. Therefore we conclude that $F$ is oriented. ∎
As yet another application of the first Koszul forms, resembling the notion of positive holomorphic line bundles in complex geometry, we now define the positivity of a flat $\mathbb{R}$-line bundle, and use it to characterize Hessian metrics of Koszul type.
Definition 2.49.
Let $(M,\nabla)$ be an affine manifold. A flat line bundle $(\mathscr{L},D)$ over $M$ is said to be a positive flat line bundle over $(M,\nabla)$ if there exists a bundle metric $\langle\cdot,\cdot\rangle$ of $\mathscr{L}\rightarrow M$ such that $\nabla\varkappa>0$, where $\varkappa$ is the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$.
Proposition 2.50.
An affine manifold admits a Hessian metric of Koszul type if and only of it admits a positive flat line bundle.
Proof.
Let $(M,\nabla)$ be an arbitrary but fixed affine manifold.
Assume that there exists a Hessian metric $h$ on $(M,\nabla)$ and a closed $1$-form $\eta\in\Omega^{1}(M)$ such that $h=\nabla\eta$. Consider an open covering $\{U_{\alpha}\}_{\alpha\in I}$ of $M$ and for each $\alpha\in I$ a smooth function $f_{\alpha}\in C^{\infty}(U_{\alpha})$, such that for each $\alpha\in I$ it holds that $\eta_{\alpha}=df_{\alpha}$, where $\eta_{\alpha}$ is the restriction of $\eta$ on $U_{\alpha}$. The flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$,
Let $\mathscr{L}\rightarrow M$ be the $\mathbb{R}$-line bundle over $M$ defined by the 1-cocycle $\{\operatorname{exp}(f_{\alpha}-f_{\beta})\}_{\alpha,\beta\in I}$ and let $D$ be the flat connection on $\mathscr{L}\rightarrow M$ defined by the connection 1-forms $\{df_{\alpha}\}_{\alpha\in I}$.
Let $\{s_{\alpha}\}_{\alpha\in I}$ be an affine local frame of $(\mathscr{L},D)$. Without loss of generality, for each $\alpha\in I$, let $U_{\alpha}$ be the domain of definition of $s_{\alpha}$. Define a bundle metric $\langle\cdot,\cdot\rangle$ on $\mathscr{L}$ by
$\langle s_{\alpha},s_{\alpha}\rangle:=\operatorname{exp}(-2f_{\alpha})$
for every $\alpha\in I$. Then, $\langle\cdot,\cdot\rangle$ is well-defined, and by construction the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is precisely $\eta$. Therefore $\nabla\eta=h>0$, and hence $(\mathscr{L},D)$ is positive.
Conversely, if $(M,\nabla)$ admits a positive flat line bundle $(\mathscr{L},D)$, then by definition there exists a bundle metric $\langle\cdot,\cdot\rangle$ such that $h:=\nabla\varkappa>0$ is a Hessian metric of Koszul type, where $\varkappa$ is the first Koszul form of the Hessian line bundle $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$. ∎
3 Interaction with Classical Geometric Structures
3.1 Digression on Complex Geometry
This digression demonstrates the utility of our framework beyond affine geometry. By lifting the Hessian structure to the tangent bundle, we can leverage powerful theorems from complex analysis.
We prove that every positive flat line bundles over an affine manifold $(M,\nabla)$ can be lifted to a positive holomorphic line bundles over the Sasaki lifting of $(M,\nabla)$, and that the Sasaki lifting of a complete Hessian manifold is Stein, which finally results in an analytic embedding theorem for complete Hessian manifolds.
Let $(M,\nabla)$ be an affine manifold, and let $\pi\colon TM\rightarrow M$ be the tangent bundle of $M$. Denote by $\mathcal{A}$ the sheaf of affine functions on $(M,\nabla)$.
Consider any coordinate map
$U\longrightarrow\mathbb{R}^{n},\quad p\longmapsto(x^{1}(p),\dots,x^{n}(p))$
of $M$ such that $x^{1},\dots,x^{n}\in\mathcal{A}(U)$ are affine functions, and that $TM\rightarrow M$ is trivialized over $U$.
Then, it is readily seen that
$\pi^{-1}(U)\longrightarrow\mathbb{R}^{2n},\quad(p,v^{1}\partial/\partial x^{1}+\cdots+v^{n}\partial/\partial x^{n})\longmapsto(x^{1}(p),\dots,x^{n}(p);v^{1},\dots,v^{n})$
is a coordinate map of $TM$.
Now, the canonical almost complex structure $J_{\nabla}$ on the tangent bundle $T(TM)\rightarrow TM$ is defined by
$J_{\nabla}(\partial/\partial x^{i}):=-\partial/\partial v^{i},\quad J_{\nabla}(\partial/\partial v^{i}):=\partial/\partial x^{i}$
for all $i=1,\dots,n$. By the classical construction in [24], the linear operator $J_{\nabla}$ is a well-defined integrable complex structure, i.e. $(TM,J_{\nabla})$ is a complex manifold. The complex manifold $(TM,J_{\nabla})$ is said to be the Sasaki lifting of $(M,\nabla)$, and the holomorphic coordinates
$z^{1}:=x^{1}+\sqrt{-1}v^{1},\dots,z^{n}:=x^{n}+\sqrt{-1}v^{n}$
of $(TM,J_{\nabla})$ are termed the canonical complexifications of the affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$.
Moreover, for each flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$, it is readily seen that the anti-holomorphic part $\bar{\partial}_{D}$ of the $\mathbb{C}$-linear extension of the connection $D$ is a Dolbeault operator on the pullback bundle $\mathscr{L}_{\mathbb{C}}:=\pi^{*}(\mathscr{L}\otimes\mathbb{C})$.
We thus arrived at the following definition:
Definition 3.1.
For an affine manifold $(M,\nabla)$, the complex manifold $(TM,J_{\nabla})$ constructed above is termed the Sasaki lifting of the affine manifold $(M,\nabla)$, and moreover, for a flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$, the holomorphic line bundle $(\mathscr{L}_{\mathbb{C}},\bar{\partial}_{D})$ over $(TM,J_{\nabla})$ constructed above is termed the holomorphic pullback of $(\mathscr{L},D)$ along $\nabla$, where a holomorphic line bundle here is equivalently defined as a pair consists of a smooth $\mathbb{C}$-line bundle and a Dolbeault operator on it.
For the complex geometric aspect of affine manifolds, the following result is of fundamental importance.
Lemma 3.2.
Let $(M,\nabla)$ be an affine manifold, and $(TM,J_{\nabla})$ its Sasaki lifting. Suppose that $(\mathscr{L},D)$ is a positive flat line bundle over $(M,\nabla)$. Then the holomorphic pullback $(\mathscr{L}_{\mathbb{C}},\bar{\partial}_{D})$ of $(\mathscr{L},D)$ along $\nabla$ is a positive holomorphic line bundle over the complex manifold $(TM,J_{\nabla})$.
Proof.
By the positivity of $(\mathscr{L},D)$, there exists a bundle metric $\langle\cdot,\cdot\rangle$ of $\mathscr{L}\rightarrow M$ such that the first Koszul form $\varkappa$ of the Hessian line bundle $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ satisfies $\nabla\varkappa>0$.
Take any arbitrary point $p\in M$. Since the first Koszul form $\varkappa$ of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is closed, by Poincaré’s lemma there exists an open neighborhood $U$ of $p$ in $M$ and a smooth function $f\in C^{\infty}(U)$ such that $\varkappa|_{U}=df$. Then, by the very definition of the first Koszul form of a Hessian line bundle, there exists a non-vanishing parallel local section $s$ of $\mathscr{L}\rightarrow M$ defined on $U$, such that $\langle s,s\rangle=\operatorname{exp}(-2f)$.
Denote by $\pi$ the bundle map of the tangent bundle $TM\rightarrow M$. We equip $(\mathscr{L}_{\mathbb{C}},\bar{\partial}_{D})$ with the pullback bundle metric $\langle\cdot,\cdot\rangle_{\mathbb{C}}:=\pi^{*}\langle\cdot,\cdot\rangle$. Then the curvature 2-form $\Theta$ of Chern connection of the Hermitian line bundle $((\mathscr{L}_{\mathbb{C}},\bar{\partial}_{D}),\langle\cdot,\cdot\rangle_{\mathbb{C}})$ is given by
$2\pi\sqrt{-1}\Theta=\partial\bar{\partial}\operatorname{log}(\langle\pi^{*}s,\pi^{*}s\rangle_{\mathbb{C}})=2\partial\bar{\partial}(\pi^{*}f).$
Now it remains to prove that $\pi^{*}f$ is a plurisubharmonic function. But this follows immediately from the assumption that $\nabla(df)=\nabla\varkappa|_{U}>0$. ∎
After the notion of pullback of line bundles being successfully defined, we also want to discover a proper way to lift smooth functions from the base. Instead of naively pullback smooth functions on an affine manifold to its tangent bundle via the bundle projection, there exists a much more purposeful method of lifting smooth functions.
Proposition 3.3.
Let $(M,\nabla)$ be an affine manifold, and $(TM,J_{\nabla})$ its Sasaki lifting. Let $f\colon M\rightarrow\mathbb{R}$ and $F\colon TM\rightarrow\mathbb{C}$ be smooth functions satisfying
$F(x^{1},\dots,x^{n};v^{1},\dots,v^{n})=f(x^{1},\dots,x^{n})+\sqrt{-1}\sum_{i=1}^{n}\frac{\partial f}{\partial x^{i}}(x^{1},\dots,x^{n})v^{i}$
where here
$z^{1}=x^{1}+\sqrt{-1}v^{1},\dots,z^{n}=x^{n}+\sqrt{-1}v^{n}$
are the canonical complexifications of some affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$. Then $f\in C^{\infty}(M)$ is affine if and only of $F\in C^{\infty}(TM)\otimes\mathbb{C}$ is holomorphic.
Proof.
Apply Cauchy-Riemann equation to the specific complex structure $J_{\nabla}$. ∎
Proposition 3.3 then motivates the following definition for general smooth functions.
Definition 3.4.
Let $(M,\nabla)$ be an affine manifold, and let $(TM,J_{\nabla})$ be its Sasaki lifting. Let
$z^{1}=x^{1}+\sqrt{-1}v^{1},\dots,z^{n}=x^{n}+\sqrt{-1}v^{n}$
be the canonical complexifications of some affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$. For any smooth function $f\in C^{\infty}(M)$ on $M$, The complex-valued smooth function $F\in C^{\infty}(TM)\otimes\mathbb{C}$ defined by
$F(x^{1},\dots,x^{n};v^{1},\dots,v^{n}):=f(x^{1},\dots,x^{n})+\sqrt{-1}\sum_{i=1}^{n}\frac{\partial f}{\partial x^{i}}(x^{1},\dots,x^{n})v^{i}$
is termed the canonical complexification of $f$ along $\nabla$.
Note that, the notion of canonical complexifications of affine coordinates is consistent with the definition of canonical complexifications of smooth functions, which also justifies the usage of terminology.
Although a tangent bundle of positive dimension is never compact, we may hope that the Sasaki liftings are Stein. Fortunately, for the the Sasaki liftings of complete Hessian manifolds, this is indeed true.
Proposition 3.5.
Let $((M,\nabla),h)$ be a complete Hessian manifold, and $(TM,J_{\nabla})$ its Sasaki lifting. Then $(TM,J_{\nabla})$ is Stein.
Proof.
Let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. By considering the Sasaki lifting of $(\tilde{M},\pi^{*}\nabla)$, it is readily seen that there exists a holomorphic covering map $T\tilde{M}\rightarrow TM$. Therefore it suffices to show that $T\tilde{M}$ is Stein.
But, by Theorem 2.13, $T\tilde{M}$ is a convex tube domain, and is therefore pesudoconvex by complex analysis of several variables. ∎
The following totally real embedding theorem for complete Hessian manifolds is a direct application of Proposition 3.5.
Theorem 3.6.
Let $((M,\nabla),h)$ be an $n$-dimensional complete Hessian manifold. Then there exists a proper analytic embedding of $M$ into $\mathbb{R}^{d}$ for any integer $d>(3n+1)/2$.
Proof.
By Proposition 3.5, the Sasaki lifting $(TM,J_{\nabla})$ of $(M,\nabla)$ is Stein. By Stein’s embedding theorem [25], there exists a proper holomorphic embedding $\Phi$ of $(TM,J_{\nabla})$ into $\mathbb{C}^{d}$ for any integer $d>(3n+1)/2$. Now, it suffices to restrict the real part $\operatorname{Re}(\Phi)$ of $\Phi$ to the totally real submanifold $M$ of $(TM,J_{\nabla})$. ∎
Remark 3.7.
Recall that Whitney’s embedding theorem states that all real analytic $n$-manifolds can be analytically embedded into $\mathbb{R}^{2n}$. Theorem 3.6 suggests a more efficient bound on analytical embedding dimension for complete Hessian manifolds.
3.2 Legendre Transformation
@@
Let $M$ be a differentiable manifold. An atlas $\mathscr{A}=\left\{\left(U_{\alpha},\varphi_{\alpha}\right)\mid\alpha\in I\right\}$ of $M$ is said to be a radiant affine atlas, if for any $\alpha,\beta\in I$ there exist $A\in\operatorname{GL}_{n}(\mathbb{R})$ such that
@@
By Theorem 2.13, we may identify $\Omega:=\Phi(\tilde{M})=\tilde{M}$. An Euler vector field $H$ of $(M,\nabla)$ lifts to a standard Euler vector field
@@
Corollary 3.17.
Let $(M,\nabla)$ be a compact radiant manifold. Then the canonical line bundle $K_{\nabla}$ of $(M,\nabla)$ is non-trivial.
Proof.
Combine Theorem 3.15 with Corollary 2.37. ∎
@@
We shall also recall the core concept of information geometry, that is, the notion of dual connections.
Definition 3.19.
$Zg(X,Y)=g\left(D_{Z}X,Y\right)+g\left(X,D_{Z}^{*}Y\right)$
holds for all smooth vector fields $X,Y,Z$ on $M$.
Theorem 3.20.
$\bar{\nabla}:=\frac{1}{2}\left(\nabla^{*}+\nabla\right)$
is the Levi-Civita connection of $(M,g)$.
Proof.
See [15] for a proof of this statement. ∎
Lemma 3.21.
Proof.
Take arbitrary vector fields $X,Y$ on $M$. Then
$(\nabla^{*}_{X}\theta)(Y)=X(\theta(Y))-\theta(\nabla^{*}_{X}Y).$
Now, since $\theta(Y)=g(\theta^{\sharp},Y)$, we have
Moreover, since
substituting this into the previous expression, we obtain
@@
Now, in view of Corollary 3.24, for a radiant Hessian manifold $((M,\nabla),h)$ of Koszul type, it is natural to ask for an explicit expression for the $1$-form potential for its dual $((M,\nabla^{*}),h)$. We will now give an satisfactory answer to this question.
@@
Let $((M,\nabla),h)$ be a radiant Hessian manifold of Koszul type, with Euler vector field $H$ and $1$-form potential $\eta$. The differential form $\eta^{*}:=-\eta+\mathcal{L}_{H}\eta$ is said to be the Legendre transform of $\eta$ along direction $H$.

### v6

@@ 2.4 Applications of Koszul Forms
Since $\mathrm{H}^{1}(M;\mathbb{R})=0$, by Proposition 2.29 we have that the Picard group of $M$ is isomorphic to $\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})$, and hence is 2-torsion. In particular, the canonical line bundle $K$ of $(M,\nabla)$ satisfies $K^{\otimes 2}=0$. Since $M$ is orientable, the $\mathbb{R}$-line bundle $\bigwedge^{n}(T^{*}M)$ is trivial. Therefore, by virtue of Proposition 2.31, the canonical line bundle of $M$ is trivial as a flat $\mathbb{R}$-line bundle. Corollary 2.33 then implies that $M$ admits a volume form $\Omega$ such that $\nabla\Omega=0$. The desired result follows immediately from Theorem 2.40. ∎
3 Essential Tools from Analysis
3.1 Radiant Hessian manifolds
@@
Let $M$ be a differentiable manifold. An atlas
of $M$ is said to be a radiant affine atlas, if for any $\alpha,\beta\in I$ there exist $A\in\operatorname{GL}_{n}(\mathbb{R})$ such that
@@
By Theorem 2.12, there exists a convex domain $\Omega$ in $\mathbb{R}^{n}$ such that $\Omega:=\Phi(\tilde{M})=\tilde{M}$. An Euler vector field $H$ of $(M,\nabla)$ lifts to a standard Euler vector field
@@
3.2 Legendre Transformation
@@
Now, in view of Corollary 3.13, for a radiant Hessian manifold $((M,\nabla),h)$ of Koszul type, it is natural to ask for an explicit expression for the $1$-form potential for its dual $((M,\nabla^{*}),h)$.
@@
Let $((M,\nabla),h)$ be a radiant Hessian manifold of Koszul type, with Euler vector field $H$ and $1$-form potential $\eta$. The differential form
$\eta^{*}:=-\eta+\mathcal{L}_{H}\eta$
is said to be the Legendre transform of $\eta$ along direction $H$.
@@
Being both radiant and of Koszul type may seem to be a strong assumption on Hessian manifolds. However, when the underlying manifold is compact, a subtle observation of J.Vey shows that the radiant condition is redundant.
Definition 3.16.
Proposition 3.17 (Vey,1970[26]).
Let $\Omega$ be a regular convex domain in $\mathbb{R}^{n}$. Suppose there exists a discrete subgroup $G$ of $\operatorname{Aff}(\Omega)$ acting freely and proper discontinuously on $\Omega$ so that $\Omega/G$ is compact. Then $\Omega$ is a cone.
We also recall a hyperbolicity condition in the context of affine differential geometry.
Definition 3.18.
Let $(M,\nabla)$ be an affine manifold, and let $\pi\colon\tilde{M}\rightarrow M$ be the universal covering of $M$. The affine manifold $(M,\nabla)$ is said to be hyperbolic, if the affine development $\Phi\colon\tilde{M}\rightarrow\mathbb{R}^{n}$ of $(\tilde{M},\pi^{*}\nabla)$ is a diffeomorphism onto its image, and its image is a regular convex domain in $\mathbb{R}^{n}$.
We observe that the notion of affine hyperbolicity is closed related to that of a Hessian metric of Koszul type.
Theorem 3.19.
Proof.
This theorem is proved in [23]. ∎
The following Corollaries are then natural and facile.
Corollary 3.20.
A compact Hessian manifold is radiant if and only if it is of Koszul type.
Proof.
By Theorem 3.19 and Proposition 3.17, a compact Hessian manifold of Koszul type is automatically radiant.
Let $((M,\nabla),h)$ be a compact radiant Hessian manifold. Then by Proposition 3.12, we have that $((M,\nabla^{*}),h)$ is a compact Hessian manifold of Koszul type, and hence is radiant. Now, since $\nabla^{**}=\nabla$, the Hessian manifold $((M,\nabla),h)$ is also of Koszul type by Corollary 3.13. ∎
Corollary 3.21.
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type. Then the $1$-form potential of $((M,\nabla),h)$ is unique.
Proof.
Recall Corollary 3.20, and then apply Corollary 3.9 and Propositon 3.11. ∎
Corollary 3.22.
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type. Then the Riemannian manifold $(M,h)$ is not flat.
Proof.
This assertion follows from Corollary 3.20 and Propositon 3.10. ∎